the end of a ladder is sitting on the ground 6 feet from a wall and as it leans against a wall the top of the ladder hits a spot 8 feet up the wall. how long is the ladder?
Answer by mananth(16949)

The ladder the ground and the wall form a right triangle

The ladder is the hypotenuse of the triangle
the height at which it touches the wall is leg1
The distance from the wall to the base of the ladder is leg2

Pythagoras theorem
Hypotenuse ^2 = leg1^2+leg2^2

leg1= 8 ft
leg2= 6 ft
Hypotenuse^2= 8^2+6^2
Hypotenuse^2= 64 + 36
Hypotenuse^2= 100
Hypotenuse = sqrt%09100
Hypotenuse = 10 ft

The height of ladder = 10 ft
m.ananth@hotmail.ca